A rubber ball is shot straight up from the ground at 4.90 m/s. A second rubber ball is simultaneously dropped from rest 1.61 m above the ground. How high above the ground are the balls when they collide?


1
Expert's answer
2020-09-21T08:30:26-0400
"y_1=h-0.5gt^2,y_2=vt-0.5gt^2\\y_1=y_2\to h-0.5gt^2=vt-0.5gt^2"

"t=\frac{h}{v}=\frac{1.61}{4.9}=0.329\ s"

"y=(4.9)(0.329)-0.5(9.8)(0.329)^2=1.08\ m"